OpenStudy (anonymous):

How do i figure out this problem - 5/9p = 10 help please!!!

OpenStudy (texaschic101):

-5/9p = 10 --- multiply both sides by 9 -5p = 90 -- divide both sides by -5 p = -90/5 p = - 18

OpenStudy (anonymous):

Thanx to both you guys texaschic101 and Sierra